A hotplate has an area of 0.15m^2 and temperature of 300 C the room temperature is 40 C how much net power is transmitted to room in every second? e=0.80. what are the radiated wave length from hot plate and room?
Answer #1

Solution) A = 0.15 m^2

T1 = 300°C = 300 + 273 = 573 K

T2 = 40°C = 40 + 273 = 313 K

e = 0.80

Power , P = ?

P = Ae(sigma)( (T1)^(4) - (T2)^(4))

Sigma = 5.67×10^(-8) W/(m^2)(K^4)

P = 0.15×0.80×5.67×10^(-8)((573)^4 - (313)^4)

P = 668.16 W

Wavelength , Lambda = ? For hot plate T1 = 300°C

T1 = 300 + 273 = 573 K

From weins law

Wavelength, Lambda = (2.898×10^(-3))/(T)

Lambda = (2.898×10^(-3))/(573)

Lambda = 5.05×10^(-6) m

Wavelength, Lambda = ? For room T2 = 40°C = 313 K

Lambda = (2.898×10^(-3))/(313)

Lambda = 9.25×10^(-6) m